Hydrogen atom in its ground is excited by means of monochromatic radiation of wavelength 975 Å. How many different lines are possible in the resulting spectrum ? Calculate the longest wavelength amongst them. You may assume the ionization energy for hydrogen atom as 13.6 eV.

Grade:upto college level

2 Answers

Rinkoo Gupta
9 years ago
Four Lines are possible in the resulting spectrum.
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
Deepak Patra
9 years ago
. \frac{238}{92}U ⟶ \frac{234}{90}X + \frac{4}{2}He \frac{234}{91}X ⟶ \frac{234}{90}Y + \frac{0}{-1}e
(i) Atomic number = 91
(ii) Mass number = 234